OpenStudy (lolipop_drum312):

What is the justification for each step in the solution of the equation 3(2π‘₯+7)βˆ’4=53βˆ’3π‘₯?

OpenStudy (lolipop_drum312):

3(2π‘₯+7)βˆ’4=53βˆ’3π‘₯ _______________ 6π‘₯ + 21 – 4 = 53 – 3π‘₯ ______________ 6π‘₯ + 17 = 53 – 3π‘₯ __________________ 9π‘₯ + 17 = 53 ________________ 9π‘₯ = 36 _______________ π‘₯ = 4 _______________

OpenStudy (lolipop_drum312):

i think for the first one is distributive

OpenStudy (lolipop_drum312):

@welshfella

OpenStudy (welshfella):

yea but I'm not very good at these justifications though Sorry

OpenStudy (lolipop_drum312):

ok

OpenStudy (lolipop_drum312):

my theory: Subtract property of equality Addition property of equality Subtract property of equality Division property of equality Distributive property Given

OpenStudy (lolipop_drum312):

idk tho

OpenStudy (lolipop_drum312):

@Kevin

OpenStudy (kevin):

I'm not good at justification either. Maybe I'll answer your question later. Sorry

OpenStudy (lolipop_drum312):

ok

OpenStudy (welshfella):

we just learned the rules but we were never given labels for them.

OpenStudy (lolipop_drum312):

the labels for them?

OpenStudy (welshfella):

the 'Addition Property of Equality' and so on.

OpenStudy (welshfella):

I think that one is a = b therefore a + d = b + d

OpenStudy (welshfella):

- seems common sense to me. Why label it?

OpenStudy (welshfella):

anyway - that's me....

OpenStudy (lolipop_drum312):

@NerdyChick16

OpenStudy (nerdychick16):

@awesomedude2078 @GracyGirl @GracyGirl @desmarie @LazyBoy @jango_IN_DTOWN

OpenStudy (nerdychick16):

@Shy_Boy

OpenStudy (gracygirl):

Wait so like a two column proof?

OpenStudy (lolipop_drum312):

it like a justifications thing naming what property they are

OpenStudy (lolipop_drum312):

examples up their ^

OpenStudy (gracygirl):

3(2π‘₯+7)βˆ’4=53βˆ’3π‘₯ -Given 6π‘₯ + 21 – 4 = 53 – 3π‘₯ -Distributive Property 6π‘₯ + 17 = 53 – 3π‘₯ -Subtraction Property of Equality 9π‘₯ + 17 = 53 -Subtraction Property of Equality 9π‘₯ = 36 -Addition Property Of Equality π‘₯ = 4 -Division Property of Equality I believe that is correct.

OpenStudy (lolipop_drum312):

ok thank you

OpenStudy (gracygirl):

Actually.... 3(2π‘₯+7)βˆ’4=53βˆ’3π‘₯ -Given 6π‘₯ + 21 – 4 = 53 – 3π‘₯ -Distributive Property 6π‘₯ + 17 = 53 – 3π‘₯ -Addition Property of Equality 9π‘₯ + 17 = 53 -Subtraction Property of Equality 9π‘₯ = 36 -Subtraction Property of Equality π‘₯ = 4 -Division Property of Equality There that's correct.

OpenStudy (lolipop_drum312):

oooh ok

OpenStudy (gracygirl):

:) Anytime